OpenStudy (anonymous):

The area of a square lawn is 196 yd2. What is the length of each side of the lawn?

OpenStudy (anonymous):

Formula of Area of square = (side)^2 here given that area = 196 so 196 yd^2 = (side)^2 we have to take square root of both sides. so we get 14 = side. each side is equal in the case of square. so each side = 14 yd.